# Asymptotic theory of in-context learning by linear attention

### v2

@@ Acknowledgments
$e(\Gamma^{\ast})\simeq e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda),$ (SI.5.28)
@@
$e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ $\coloneqq\left(\frac{1+\rho}{\alpha}+1\right)\left(1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)-c_{e}\left[\mathcal{M}_{\kappa}(\nu)+\xi\mathcal{M}^{\prime}_{\kappa}(\nu)\right]\right)$ (SI.5.29)
@@
Recall the definition of the asymptotic equivalence notation $\simeq$” introduced in Section SI.1. The characterization given in (SI.5.28) implies that, as $d\to\infty$, the generalization error $e(\Gamma^{\ast})$ converges almost surely to the deterministic quantity $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$.
To derive (SI.5.28), our starting point is the estimate
$e(\Gamma^{\ast})\simeq\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}B_{\mathrm{test}}(\Gamma^{\ast})^{\top}\right)-\frac{2}{d}\operatorname{tr}\left(\Gamma^{\ast}A_{\mathrm{test}}^{\top}\right)+1+\rho,$ (SI.5.30)
which follows from Corollary 1 and the discussions in Remark 2. We consider the ICL task here, and thus $A_{\mathrm{test}}$ and $B_{\mathrm{test}}$ are given in (SI.5.1). The asymptotic limits of the first two terms on the right-hand side of the above equation can be obtained by the characterizations given in Result 4.
Using (SI.5.3) and the expressions in (SI.5.21) and (SI.5.1), we have
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})$ $\simeq\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{test}}^{\top}\right)$ (SI.5.31)
$=1-\frac{\nu}{d}\operatorname{tr}F_{R}(\nu)-a^{\ast}(1+\rho)^{2}\nu\frac{\lVert F_{R}(\nu)b_{\mathrm{tr}}\rVert^{2}}{d}$ (SI.5.32)
$\simeq 1-\nu\mathcal{M}_{\kappa}(\nu),$ (SI.5.33)
where $\nu$ is the constant defined in (SI.5.8). To reach the last step, we have used the estimate given in (SI.8.19).
Next, we use (SI.5.4) to characterize the first term on the right-hand side of (SI.5.30). From the formulas in (SI.5.21) and (SI.5.1), we can check that
$\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top}\right)$ $\simeq\left(\frac{1+\rho}{\alpha}+1\right)\frac{1}{d}\operatorname{tr}\left(I-\nu F(\nu)\right)^{2}$ (SI.5.34)
$\simeq\left(\frac{1+\rho}{\alpha}+1\right)\left(1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right),$ (SI.5.35)
where the second step follows from (SI.8.19) and (SI.8.20). From (SI.8.7),
$\frac{1}{d}\operatorname{tr}(B_{\mathrm{test}}(E_{\mathrm{tr}}+\xi I)^{-1})\simeq\left(\frac{1+\rho}{\alpha}+1\right)\frac{1}{d}\operatorname{tr}F_{R}(\nu)\simeq\left(\frac{1+\rho}{\alpha}+1\right)\mathcal{M}_{\kappa}(\nu).$ (SI.5.36)
@@
$\frac{1}{d}\operatorname{tr}(B_{\mathrm{test}}(E_{\mathrm{tr}}+\xi I)^{-2})\simeq\left(\frac{1+\rho}{\alpha}+1\right)\frac{1}{d}\operatorname{tr}F^{2}_{R}(\nu)\simeq-\left(\frac{1+\rho}{\alpha}+1\right)\mathcal{M}^{\prime}_{\kappa}(\nu).$ (SI.5.37)
Substituting (SI.5.35), (SI.5.36), and (SI.5.37) into (SI.5.4) gives us
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B(\Gamma^{\ast})^{\top})\simeq\left(\frac{1+\rho}{\alpha}+1\right)\left(1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)-c_{e}\left[\mathcal{M}_{\kappa}(\nu)+\xi\mathcal{M}^{\prime}_{\kappa}(\nu)\right]\right),$ (SI.5.38)
where $c_{e}$ is the constant given in (SI.5.7). Combining (SI.5.33), (SI.5.38), and (SI.5.30), we are done.
@@
$q^{\ast}\coloneqq\frac{1+\rho}{\alpha},\qquad m^{\ast}\coloneqq\mathcal{M}_{\kappa}\left({q^{\ast}}\right),\qquad\text{and}\qquad\mu^{\ast}\coloneqq q^{\ast}\mathcal{M}_{\kappa/\tau}(q^{\ast}),$ (SI.5.39)
where $\mathcal{M}_{\kappa}(x)$ is the function defined in (SI.8.3). Then
$e^{\mathrm{ICL}}_{\mathrm{ridgeless}}\coloneqq\lim_{\lambda\to 0^{+}}e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ (SI.5.40)
=$=\begin{cases}{\frac{\tau(1+q^{\ast})}{1-\tau}\left[1-\tau(1-\mu^{\ast})^{2}+\mu^{\ast}(\rho/q^{\ast}-1)\right]}{-2\tau(1-\mu^{\ast})+(1+\rho)}&\quad\tau<1\\ \left(q^{\ast}+1\right)\left(1-2q^{\ast}m^{\ast}-(q^{\ast})^{2}\mathcal{M}^{\prime}_{\kappa}(q^{\ast})+\frac{(\rho+q^{\ast}-(q^{\ast})^{2}m^{\ast})m^{\ast}}{\tau-1}\right)-2(1-q^{\ast}m^{\ast})+(1+\rho)&\quad\tau>1\end{cases},$
@@
We start with the case of $\tau<1$. Examining the self-consistent equation in (SI.5.6), we can see that the parameter $\xi$ tends to a nonzero constant, denoted by $\xi^{\ast}$, as $\lambda\to 0^{+}$. It follows that the original equation in (SI.5.6) reduces to
${\xi^{\ast}}\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi^{\ast}\right)=1-\tau.$ (SI.5.41)
@@
$\mu^{\ast}\coloneqq\frac{(1-\tau)(1+\rho)}{\alpha\tau\xi^{\ast}}.$ (SI.5.42)
By combining (SI.5.41) and the characterization in (SI.8.4), we can directly solve for $\mu$ and get $\mu^{\ast}=q^{\ast}\mathcal{M}_{\kappa/\tau}(q^{\ast})$ as given in (SI.5.39). The characterization in (SI.5.40) (for the case of $\tau<1$) then directly follows from (SI.5.33), (SI.5.38), and (4) after some lengthy calculations.
Next, we consider the case of $\tau>1$. It is straightforward to verify from (SI.5.6) that
$\xi=\frac{\tau}{\tau-1}\lambda+\mathcal{O}(\lambda^{2}).$ (SI.5.43)
@@
$\lim_{\lambda\to 0^{+}}\nu=\lim_{\lambda\to 0^{+}}\left(\frac{1+\rho}{\alpha}+\xi\right)=q^{\ast}\quad\text{and}\quad\lim_{\lambda\to 0^{+}}\mathcal{M}_{\kappa}(\nu)=m^{\ast}.$ (SI.5.44)
Substituting these estimates into (SI.5.33), (SI.5.38), and (4), we then reach the characterizations in (SI.5.40) for the case of $\tau>1$.
SI.5.3 The Generalization Error of In-Distribution Generalization
@@
$e(\Gamma^{\ast})\simeq e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)\coloneqq\tau\frac{\rho+\nu-\nu^{2}\mathcal{M}_{\kappa}(\nu)-\xi\left[1-{2\nu}\mathcal{M}_{\kappa}(\nu)-{\nu^{2}}\mathcal{M}^{\prime}_{\kappa}(\nu)\right]}{\tau-\left[1-{2\xi}\mathcal{M}_{\kappa}(\nu)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right]},$ (SI.5.45)
where $\xi$ the unique positive solution to the self-consistent equation (SI.5.6) and $\nu$ is the constant given in (SI.5.8).
Similar to our derivation of Result 5, we only need to use (SI.5.3) and (SI.5.4) to characterize the asymptotic limits of the first and second terms on the right-hand side of (SI.5.30). Note that, for the IDG task, $A_{\mathrm{test}}=A_{\mathrm{tr}}$. It follows from (SI.5.3) and (SI.5.24) that
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})\simeq 1-\nu+\nu^{2}\mathcal{M}_{\kappa}(\nu).$ (SI.5.46)
Similarly, since $B_{\mathrm{test}}=E_{\mathrm{tr}}$, we can verify from (SI.5.5) that
$\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top}\right)$ $=\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{tr}}^{\top})-\frac{\xi}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})$ (SI.5.47)
$\simeq 1-\nu+\nu^{2}\mathcal{M}_{\kappa}(\nu)-\xi\left[1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right],$ (SI.5.48)
where the second step follows from (SI.5.24) and (SI.5.25). Moreover,
$\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}\left[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}\right]\right)=1-2\xi\mathcal{M}_{\kappa}(\nu)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}(\nu).$ (SI.5.49)
Substituting (SI.5.47) and (SI.5.49) into (SI.5.4), we have
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B(\Gamma^{\ast})^{\top})$ (SI.5.50)
@@
The final result in (SI.5.45) then follows from combining the above expression with (SI.5.46) and (SI.5.30).
@@
Let $q^{\ast}$, $m^{\ast}$, and $\mu^{\ast}$ be the scalars defined in (SI.5.39). We have
$e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ $\coloneqq\lim_{\lambda\to 0^{+}}e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)=\begin{cases}\frac{\tau}{1-\tau}\left(\frac{\rho+q^{\ast}-2q^{\ast}(1-\tau)({q^{\ast}}/{\xi^{\ast}}+1)}{1-p^{\ast}(1-\tau)}+\frac{\tau\mu^{\ast}(q^{\ast}+\xi^{\ast})^{2}}{q^{\ast}}\right)&\qquad\tau<1\\ \frac{\tau}{\tau-1}[\rho+q^{\ast}(1-q^{\ast}m^{\ast})]&\qquad\tau>1\end{cases},$ (SI.5.51)
@@
The derivation of this result closely follows that of Result 6. We analyze the cases of $\tau<1$ and $\tau>1$ separately. For $\tau<1$, the equation in (SI.5.6) simplifies to (SI.5.41) as $\lambda\to 0^{+}$. For $\tau>1$, $\xi$ approaches zero as $\lambda\to 0^{+}$. Substituting these estimates into (SI.5.45) then yields (SI.5.51) after some detailed calculations.
SI.5.4 A note on ordering of $\lambda\to 0^{+}$ and $\alpha\to\infty$ limits
@@
$H_{Z}=\begin{bmatrix}x_{\ell+1}w^{\top}&\frac{1}{d}\left(\|w\|^{2}+\rho\right)x_{\ell+1}\end{bmatrix}$ (SI.5.52)
@@
$\lim_{\lambda\to 0^{+}}$ $\lim_{\alpha\to\infty}e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)=\begin{dcases}1-\kappa+\rho+\frac{\rho\kappa^{2}}{(\tau-\kappa)(1-\kappa)}&\kappa\leq\text{min}(\tau,1)\\ 1-\tau+\rho+\rho\frac{\kappa\tau}{(\kappa-\tau)(1-\tau)}&\tau<1,\kappa>\tau\\ \rho+\frac{\rho\kappa}{(\tau-1)(\kappa-1)}&\tau>1,\kappa>1\end{dcases},$ (SI.5.53)
noting that the second and third cases match their equivalent finite cases in (35) in the manuscript. This result follows from applying the asymptotics given in prior works for the generalization error of ridgeless regression with structured covariates [38, 39, 41]. By comparison with (35), we see that switching the order of limits in $\alpha$ and $\lambda$ has mitigated the divergence for $\kappa\leq\min(\tau,1)$.
As these limits do not commute, which order is correct? Simulations with arbitrarily small numerical ridge parameter at large $\alpha$ track the (35) curve, not the (SI.5.53) curve. These numerical experiments are computed with the ridge scaling $n/d$ that we use elsewhere. Were we to scale this regularisation with some power of $\alpha$, we could remove this large $\alpha$ divergence; whether or not that choice is either theoretically or practically motivated is a question we leave open for future work.
SI.6 Task Generalization and Bayesian Estimators

### v3

@@ E.2 The Generalization Error of In-Context Learning
$e(\Gamma^{\ast})\simeq e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda),$ (174)
@@
$e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ $\coloneqq\left(\frac{1+\rho}{\alpha}+1\right)\left(1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)-c_{e}\left[\mathcal{M}_{\kappa}(\nu)+\xi\mathcal{M}^{\prime}_{\kappa}(\nu)\right]\right)$ (175)
@@
Recall the definition of the asymptotic equivalence notation $\simeq$” introduced in Section A. The characterization given in (174) implies that, as $d\to\infty$, the generalization error $e(\Gamma^{\ast})$ converges almost surely to the deterministic quantity $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$.
To derive (174), our starting point is the estimate
$e(\Gamma^{\ast})\simeq\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}B_{\mathrm{test}}(\Gamma^{\ast})^{\top}\right)-\frac{2}{d}\operatorname{tr}\left(\Gamma^{\ast}A_{\mathrm{test}}^{\top}\right)+1+\rho,$ (176)
which follows from Corollary 1 and the discussions in Remark 2. We consider the ICL task here, and thus $A_{\mathrm{test}}$ and $B_{\mathrm{test}}$ are given in (147). The asymptotic limits of the first two terms on the right-hand side of the above equation can be obtained by the characterizations given in Result 4.
Using (149) and the expressions in (167) and (147), we have
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})$ $\simeq\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{test}}^{\top}\right)$ (177)
$=1-\frac{\nu}{d}\operatorname{tr}F_{R}(\nu)-a^{\ast}(1+\rho)^{2}\nu\frac{\mathinner{\!\left\lVert F_{R}(\nu)b_{\mathrm{tr}}\right\rVert}^{2}}{d}$ (178)
$\simeq 1-\nu\mathcal{M}_{\kappa}(\nu),$ (179)
where $\nu$ is the constant defined in (154). To reach the last step, we have used the estimate given in (268).
Next, we use (150) to characterize the first term on the right-hand side of (176). From the formulas in (167) and (147), we can check that
$\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top}\right)$ $\simeq\left(\frac{1+\rho}{\alpha}+1\right)\frac{1}{d}\operatorname{tr}\left(I-\nu F(\nu)\right)^{2}$ (180)
$\simeq\left(\frac{1+\rho}{\alpha}+1\right)\left(1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right),$ (181)
where the second step follows from (268) and (269). From (256),
$\frac{1}{d}\operatorname{tr}(B_{\mathrm{test}}(E_{\mathrm{tr}}+\xi I)^{-1})\simeq\left(\frac{1+\rho}{\alpha}+1\right)\frac{1}{d}\operatorname{tr}F_{R}(\nu)\simeq\left(\frac{1+\rho}{\alpha}+1\right)\mathcal{M}_{\kappa}(\nu).$ (182)
@@
$\frac{1}{d}\operatorname{tr}(B_{\mathrm{test}}(E_{\mathrm{tr}}+\xi I)^{-2})\simeq\left(\frac{1+\rho}{\alpha}+1\right)\frac{1}{d}\operatorname{tr}F^{2}_{R}(\nu)\simeq-\left(\frac{1+\rho}{\alpha}+1\right)\mathcal{M}^{\prime}_{\kappa}(\nu).$ (183)
Substituting (181), (182), and (183) into (150) gives us
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B(\Gamma^{\ast})^{\top})\simeq\left(\frac{1+\rho}{\alpha}+1\right)\left(1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)-c_{e}\left[\mathcal{M}_{\kappa}(\nu)+\xi\mathcal{M}^{\prime}_{\kappa}(\nu)\right]\right),$ (184)
where $c_{e}$ is the constant given in (153). Combining (179), (184), and (176), we are done.
@@
$q^{\ast}\coloneqq\frac{1+\rho}{\alpha},\qquad m^{\ast}\coloneqq\mathcal{M}_{\kappa}\left({q^{\ast}}\right),\qquad\text{and}\qquad\mu^{\ast}\coloneqq q^{\ast}\mathcal{M}_{\kappa/\tau}(q^{\ast}),$ (185)
where $\mathcal{M}_{\kappa}(x)$ is the function defined in (252). Then
[…]
@@
We start with the case of $\tau<1$. Examining the self-consistent equation in (152), we can see that the parameter $\xi$ tends to a nonzero constant, denoted by $\xi^{\ast}$, as $\lambda\to 0^{+}$. It follows that the original equation in (152) reduces to
${\xi^{\ast}}\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi^{\ast}\right)=1-\tau.$ (187)
@@
$\mu^{\ast}\coloneqq\frac{(1-\tau)(1+\rho)}{\alpha\tau\xi^{\ast}}.$ (188)
By combining (187) and the characterization in (253), we can directly solve for $\mu$ and get $\mu^{\ast}=q^{\ast}\mathcal{M}_{\kappa/\tau}(q^{\ast})$ as given in (185). The characterization in (186) (for the case of $\tau<1$) then directly follows from (179), (184), and Result 4 after some lengthy calculations.
Next, we consider the case of $\tau>1$. It is straightforward to verify from (152) that
$\xi=\frac{\tau}{\tau-1}\lambda+\mathcal{O}(\lambda^{2}).$ (189)
@@
$\lim_{\lambda\to 0^{+}}\nu=\lim_{\lambda\to 0^{+}}\left(\frac{1+\rho}{\alpha}+\xi\right)=q^{\ast}\quad\text{and}\quad\lim_{\lambda\to 0^{+}}\mathcal{M}_{\kappa}(\nu)=m^{\ast}.$ (190)
Substituting these estimates into (179), (184), and (4), we then reach the characterizations in (186) for the case of $\tau>1$.
E.3 The Generalization Error of In-Distribution Generalization
@@
$e(\Gamma^{\ast})\simeq e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)\coloneqq\tau\frac{\rho+\nu-\nu^{2}\mathcal{M}_{\kappa}(\nu)-\xi\left[1-{2\nu}\mathcal{M}_{\kappa}(\nu)-{\nu^{2}}\mathcal{M}^{\prime}_{\kappa}(\nu)\right]}{\tau-\left[1-{2\xi}\mathcal{M}_{\kappa}(\nu)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right]},$ (191)
where $\xi$ the unique positive solution to the self-consistent equation (152) and $\nu$ is the constant given in (154).
Similar to our derivation of Result 5, we only need to use (149) and (150) to characterize the asymptotic limits of the first and second terms on the right-hand side of (176). Note that, for the IDG task, $A_{\mathrm{test}}=A_{\mathrm{tr}}$. It follows from (149) and (170) that
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})\simeq 1-\nu+\nu^{2}\mathcal{M}_{\kappa}(\nu).$ (192)
Similarly, since $B_{\mathrm{test}}=E_{\mathrm{tr}}$, we can verify from (151) that
$\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top}\right)$ $=\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{tr}}^{\top})-\frac{\xi}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})$ (193)
$\simeq 1-\nu+\nu^{2}\mathcal{M}_{\kappa}(\nu)-\xi\left[1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right],$ (194)
where the second step follows from (170) and (171). Moreover,
$\frac{1}{d}\operatorname{tr}\left(B_{\mathrm{test}}\left[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}\right]\right)=1-2\xi\mathcal{M}_{\kappa}(\nu)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}(\nu).$ (195)
Substituting (193) and (195) into (150), we have
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B(\Gamma^{\ast})^{\top})$ (196)
@@
The final result in (191) then follows from combining the above expression with (192) and (176).
@@
Let $q^{\ast}$, $m^{\ast}$, and $\mu^{\ast}$ be the scalars defined in (185). We have
$e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ $\coloneqq\lim_{\lambda\to 0^{+}}e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)=\begin{cases}\frac{\tau}{1-\tau}\left(\frac{\rho+q^{\ast}-2q^{\ast}(1-\tau)({q^{\ast}}/{\xi^{\ast}}+1)}{1-p^{\ast}(1-\tau)}+\frac{\tau\mu^{\ast}(q^{\ast}+\xi^{\ast})^{2}}{q^{\ast}}\right)&\qquad\tau<1\\ \frac{\tau}{\tau-1}[\rho+q^{\ast}(1-q^{\ast}m^{\ast})]&\qquad\tau>1\end{cases},$ (197)
@@
The derivation of this result closely follows that of Result 6. We analyze the cases of $\tau<1$ and $\tau>1$ separately. For $\tau<1$, the equation in (152) simplifies to (187) as $\lambda\to 0^{+}$. For $\tau>1$, $\xi$ approaches zero as $\lambda\to 0^{+}$. Substituting these estimates into (191) then yields (197) after some detailed calculations.
E.4 A note on ordering of $\lambda\to 0^{+}$ and $\alpha\to\infty$ limits
@@
$H_{Z}=\begin{bmatrix}x_{\ell+1}w^{\top}&\frac{1}{d}\left(\|w\|^{2}+\rho\right)x_{\ell+1}\end{bmatrix}$ (198)
@@
$\lim_{\lambda\to 0^{+}}$ $\lim_{\alpha\to\infty}e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)=\begin{dcases}1-\kappa+\rho+\frac{\rho\kappa^{2}}{(\tau-\kappa)(1-\kappa)}&\kappa\leq\text{min}(\tau,1)\\ 1-\tau+\rho+\rho\frac{\kappa\tau}{(\kappa-\tau)(1-\tau)}&\tau<1,\kappa>\tau\\ \rho+\frac{\rho\kappa}{(\tau-1)(\kappa-1)}&\tau>1,\kappa>1\end{dcases},$ (199)
noting that the second and third cases match their equivalent finite cases
$\lim_{\alpha\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ $=\begin{dcases}\infty&\kappa\leq\text{min}(\tau,1),\\ 1-\tau+\rho+\frac{\rho\kappa\tau}{(\kappa-\tau)(1-\tau)}&\tau<1,\kappa>\tau,\\ \rho+\frac{\rho\kappa}{(\kappa-1)(\tau-1)}&\tau>1,\kappa>1.\end{dcases}$ (200)
from the main manuscript. This result follows from applying the asymptotics given in prior works for the generalization error of ridgeless regression with structured covariates [38, 39, 41]. By comparison with (200), we see that switching the order of limits in $\alpha$ and $\lambda$ has mitigated the divergence for $\kappa\leq\min(\tau,1)$.
As these limits do not commute, which order is correct? Simulations with arbitrarily small numerical ridge parameter at large $\alpha$ track the (200) curve, not the (199) curve. These numerical experiments are computed with the ridge scaling $n/d$ that we use elsewhere. Were we to scale this regularisation with some power of $\alpha$, we could remove this large $\alpha$ divergence; whether or not that choice is either theoretically or practically motivated is a question we leave open for future work.
Appendix F Task Generalization and Bayesian Estimators